# Toward a Multi-Echelon Cyber Warfare Theory: A Meta-Game-Theoretic Paradigm for Defense and Dominance

### v1

@@ 2.2 Objectives
The goals of cyber warfare are diverse, reflecting the multifaceted nature of cyberspace conflicts. One of the primary objectives is information theft and espionage, where adversaries seek to acquire sensitive data from governments, corporations, or individuals. For example, the 2015 breach of the US Office of Personnel Management, which exposed the personal data of millions of government employees, highlights how cyber warfare can be used to gather intelligence for strategic advantage. Espionage in cyberspace is often a precursor to more aggressive actions, providing adversaries with critical knowledge to plan attacks or influence decision-making processes.
Another significant goal is the disruption of critical infrastructure. Cyber operations targeting power grids, water supplies, or transportation systems can create widespread chaos and compromise public safety. The 2015 cyberattack on Ukraine’s power grid, which left hundreds of thousands without electricity, illustrates how adversaries can leverage cyber capabilities to undermine the operational stability of a nation. Such attacks often aim to weaken a target’s resilience, demonstrating vulnerability and eroding public trust in essential services.
Cyber warfare is also used to achieve economic destabilization. Adversaries can disrupt financial institutions, manipulate stock markets, or launch ransomware attacks to cripple economic activities. The WannaCry ransomware attack in 2017, which affected hundreds of organizations worldwide, including hospitals, disrupted critical operations and caused financial losses. Such actions can destabilize national economies, weaken political systems, and provide adversaries with strategic and monetary gains.
@@ 2.3 Domains
While cyberspace serves as the primary domain for cyber warfare, its effects are far-reaching, often intersecting with the physical domains of land, sea, air, and space. Cyber actions can have tangible impacts in these areas, creating real-world consequences that extend beyond digital boundaries. For instance, cyberattacks targeting the electrical grid can disrupt power supply to entire regions, affecting households, businesses, and critical services. Similarly, satellite hacking can compromise GPS systems, affecting aviation, maritime navigation, and military operations. A notable example is the alleged 2014 cyberattack on a German steel mill, where hackers accessed industrial control systems, causing significant physical damage to a blast furnace. These intersections highlight the potential for cyber operations to inflict harm in the physical world, blurring the line between cyber and traditional warfare.
Critical infrastructure is another high-stakes target in cyber warfare, given its importance for social functioning and security. Power plants, water supply systems, transportation networks, and healthcare facilities are increasingly reliant on interconnected digital systems, making them vulnerable to cyberattacks. For example, the 2021 Colonial Pipeline ransomware attack disrupted fuel supplies across the southeastern United States, causing widespread panic and economic impact beerman2023review . Similarly, attacks on healthcare networks, such as the ransomware incident affecting Ireland’s health service in 2021, delayed medical treatments and jeopardized patient safety. These incidents underscore the critical need for robust cybersecurity measures to protect infrastructure that underpins everyday life and national security.
The human domain is also a key intersection, as cyber warfare increasingly exploits social vulnerabilities to achieve its objectives. Psychological operations (psyops), misinformation campaigns, and social engineering tactics are used to manipulate individuals or groups. For example, misinformation campaigns during elections aim to influence public opinion, undermine trust in democratic institutions, or polarize societies. In the realm of social engineering, phishing attacks trick individuals into divulging sensitive information, enabling further cyber intrusions. The 2016 US presidential election saw allegations of state-sponsored misinformation campaigns on social media platforms, demonstrating how human perceptions and behaviors can be influenced through cyber means yang2025transparent . By targeting the human domain, adversaries amplify the impact of their cyber operations, turning individuals into unwitting participants in their strategies.
@@ 2.4 Resources
AI resources are increasingly integral to modern cyber warfare, enabling automated tasks that enhance operational efficiency and scalability. Artificial intelligence can be used for real-time reconnaissance, identifying vulnerabilities in adversary networks more quickly than human operators. It is also pivotal in crafting sophisticated misinformation campaigns, such as generating deepfake videos to undermine public trust. In addition, AI-driven tools can execute cyberattacks autonomously, selecting targets or adapting strategies based on changing conditions. For example, AI-powered bots might be deployed in a DDoS attack, overwhelming targeted servers with traffic while dynamically adjusting patterns to evade countermeasures.
@@ 2.5 Operational Elements
At the highest level, strategy encompasses comprehensive planning that guides the sequencing of operations to accomplish long-term objectives. For example, a nation-state may strategically prioritize disabling an adversary’s critical infrastructure during a conflict to create chaos and weaken its operational capabilities. This type of strategic planning often includes the identification of key targets, the allocation of resources, and the coordination of multiple campaigns to ensure maximum impact and efficiency.
Operations are practical implementations of strategies that encompass both defensive and offensive efforts in the realm of cyber warfare. Defensive operations focus on protecting systems by mitigating vulnerabilities, detecting threats, and neutralizing attacks. For instance, a government agency might deploy advanced intrusion detection systems to thwart cyber intrusions targeting sensitive data. These measures are essential to maintain the integrity and security of critical infrastructure. In contrast, offensive operations involve proactive measures that aim to disrupt or disable adversary systems. An example of this would be the launch of a cyber attack on the command and control servers of an enemy to disrupt their military operations. Such actions are designed to impair the adversary’s capabilities and gain a strategic advantage.
To effectively understand and execute these operations, the concept of the kill chain has been introduced. Similar to traditional military operations, the kill chain in cyberspace involves understanding how adversaries plan and execute cyber attacks. This framework can be used to capture and model offensive operations, allowing defenders to anticipate and counteract potential threats. Similarly, a defense chain can be utilized to describe defensive operations. This concept outlines the steps taken to protect systems from cyber threats, emphasizing the importance of a structured approach to defense in the ever-evolving landscape of cyber warfare. Together, these frameworks provide valuable insight into offensive and defensive strategies in cyberspace.
@@
Within operational frameworks, information operations play a critical role. These involve tactics such as spreading misinformation, employing deception, and conducting psychological operations (psyops) to manipulate perceptions and gain tactical advantages 10685605 . For example, during a geopolitical crisis, a state might disseminate false information to create confusion and delay the adversary’s response, setting the stage for a decisive offensive cyber operation.
Attacks represent the specific engagements within cyber warfare, akin to battles in traditional conflict. These can take various forms. For example, Advanced Persistent Threats (APTs) are a class of sophisticated, stealthy intrusions that persist over extended periods, such as an APT group infiltrating a financial system to siphon funds or gather intelligence alshamrani2019survey . DDoS Attacks can overwhelm a network with traffic to render it unusable, often employed to disrupt critical services like online banking or government portals. In addition, Jamming Communications blocks or interferes with communication signals, potentially crippling military coordination or emergency response systems.
Finally, tactics involve the specific methods and techniques used to execute these attacks. These are often guided by established frameworks like MITRE ATT&CK strom2018mitre , which provides detailed documentation of adversarial Techniques, Tactics, and Procedures (TTPs). For example, a cybercriminal group might use a phishing campaign to deliver malware (technique), followed by privilege escalation (tactic), to gain control over critical systems. By studying and leveraging frameworks like MITRE ATT&CK, defenders can better anticipate and counteract these tactics, enhancing their resilience against evolving threats.
@@ 2.6 External Factors
Cyber operations are significantly shaped by external considerations, which establish the boundaries and constraints within which cyber actors operate. Among these factors, compliance with legal and regulatory frameworks is paramount. Nations, organizations, and alliances must adhere to international treaties, national laws, and organizational policies that govern the conduct of cyber operations. For example, the Tallinn Manual provides guidelines on the applicability of international law to cyber warfare, defining acceptable conduct during conflicts schmitt2017tallinn . Similarly, organizations must comply with standards such as the General Data Protection Regulation (GDPR) in the European Union, which governs the handling of personal data and restricts certain cyber activities. Non-compliance can result in legal penalties, diplomatic fallout, or reputational damage, emphasizing the need for operations that align with established norms.
Constraints also play a critical role in shaping cyber operations, encompassing technical, geopolitical, and resource limitations. Technological constraints, such as the availability of advanced tools or expertise, may restrict the scope and sophistication of an operation. For instance, a state with limited cybersecurity infrastructure might struggle to defend against or execute advanced persistent threats (APTs). Geopolitical factors, such as alliances or rivalries, influence the planning and execution of cyber campaigns. A nation might hesitate to launch a cyberattack if it risks escalating tensions with powerful adversaries or alienating allies. Resource limitations, including financial constraints or human resource shortages, can further hinder operations. An example is a small nation relying on third-party contractors for cybersecurity defense, which may leave it vulnerable to more sophisticated state-sponsored attackers.
Finally, ethics introduces an important dimension to cyber operations, dictating what actions are considered acceptable based on societal norms or organizational principles. Ethical considerations often address questions such as whether deception, misinformation, or harm to civilian systems is permissible. For instance, targeting a hospital’s network during a conflict would be deemed unethical and potentially a violation of international humanitarian law, even if the attack disrupts an adversary’s operations. On the other hand, cyber deception Zhu2025CyberDeception , such as deploying honeypots to mislead attackers, is often considered an ethical and effective defensive measure. However, ethical norms can vary widely across cultures and organizations, leading to differing interpretations of acceptable conduct. For example, some nations may see cyber espionage as a legitimate extension of traditional intelligence-gathering, while others view it as a violation of sovereignty.
@@ 2.7 Outcomes
The outcomes of cyber warfare are multifaceted and can be evaluated at various levels, depending on the objectives and the broader context of the conflict. At the most basic level, operations are assessed as a win, a loss, or a neutral, reflecting their immediate effectiveness. For example, a successful DDoS attack that temporarily disrupts an adversary’s critical infrastructure might be considered a “win” in the context of the operation, achieving its intended goal of creating disruption. Conversely, if a cyber defense system thwarts such an attack without significant impact, it would count as a “loss” for the attacker. Neutral outcomes occur when the operation fails to produce meaningful effects or when the benefits and costs balance out, leaving neither party significantly advantaged.
However, the outcomes of cyber warfare extend beyond isolated engagements, encompassing multilevel outcomes that reflect the interplay between individual operations and overarching strategic goals. In this context, success or failure in a specific operation, a “battle”, does not necessarily determine the success or failure of the larger campaign, the “war.” For instance, an Advanced Persistent Threat (APT) operation might fail to infiltrate a high-value target due to robust defenses. While this may seem like a loss, the reconnaissance conducted during the attempt could yield valuable intelligence that enables future operations to succeed, contributing to a longer-term strategic objective.
@@
It is important to recognize that outcomes in cyber warfare are not always zero-sum. Both parties can suffer losses, and the consequences can extend beyond the primary actors, resulting in significant collateral damage.
Cyber warfare can be characterized using a framework defined by key elements. Each element corresponds to a game-theoretic counterpart. Clearly defining these elements facilitates the formulation of a game, providing valuable insights for its analysis, which will be explored in the next section.
@@ 3.1 Multi-Echelon Warfare
At the policy level, cyber warfare is linked to grand strategy, sometimes referred to as high strategy, which extends beyond the immediate concerns of military and cyber operations. Grand strategy integrates economic relations, diplomatic behavior, and international collaboration to achieve long-term objectives. It involves leveraging all available instruments of national power, diplomatic, informational, military, and economic (the DIME framework) to ensure the security and prosperity of a nation or coalition. This level addresses the nation-state scale, where policies shape the operational and tactical choices that follow.
@@
At the strategy level, the focus is on high-level planning to align operations with the overarching objectives of warfare. This involves developing a sequence of defensive and offensive operations, allocating resources to maximize effectiveness, and ensuring coordination between interdependent operations. For example, if the objective is to disrupt an adversary’s critical infrastructure, the strategy could involve a combination of information operations to create misinformation and psychological pressure, followed by a cyber offensive targeting the operating technology systems of a power grid. In this context, information operations, such as the spreading of false data on the stability of the grid, could create confusion and panic, thus softening defenses and aiding the subsequent offensive operation. Resource allocation plays a pivotal role at this level, ensuring that skilled personnel, zero-day vulnerabilities, and AI-powered tools are optimally deployed in planned operations.
@@
At the operational level, the focus narrows to planning specific operations that align with strategic objectives. This involves designing a sequence of attacks that, when executed, achieve the intended outcome of an operation. For instance, in an offensive operation aimed at disabling a transportation system, the plan might involve multiple steps: first, a phishing campaign to gain initial access to the transportation agency’s network; next, a ransomware attack to encrypt critical operational data and disable system functionality; and finally, a DoS attack to prevent recovery efforts. These attacks are not isolated, but are composed in a way that each attack amplifies the impact of the others, ensuring that the operation achieves its intended goal. The success of an operation depends on the ability to identify vulnerabilities, sequence attacks effectively, and adapt to the adversary’s countermeasures.
@@
The tactic level focuses on the detailed execution of individual attacks, determining the specific techniques, tactics, and procedures (TTPs) required to carry them out successfully. For example, in a phishing attack, tactics might include crafting highly personalized emails that mimic trusted sources, leveraging psychological manipulation to increase the likelihood of the target clicking on malicious links. In a DoS attack, tactics might involve deploying botnets to overwhelm servers with traffic while simultaneously masking the source of the attack to delay mitigation. The MITRE ATT&CK framework is often used at this level to plan and refine tactics, ensuring that each attack is optimized for maximum impact.
@@
The three levels of cyber warfare are highly interconnected. Strategic planning provides the overarching framework that dictates the goals and allocation of resources for operations. The operational level translates these strategic goals into actionable plans composed of multiple coordinated attacks. The tactical level ensures that each individual attack within an operation is meticulously planned and executed to effectively contribute to the overall success. For example, a campaign to disrupt an adversary’s communication infrastructure might strategically plan to exploit interdependencies between networks and users, operationally design a series of cascading attacks on routers and satellites, and tactically deploy precision malware to exploit specific vulnerabilities in the routers.

### v2

@@ 2.2 Objectives
The goals of cyber warfare are diverse, reflecting the multifaceted nature of cyberspace conflicts. One of the primary objectives is information theft and espionage, where adversaries seek to acquire sensitive data from governments, corporations, or individuals. For example, the 2015 breach of the US Office of Personnel Management, which exposed the personal data of millions of government employees, highlights how cyber warfare can be used to gather intelligence for strategic advantage The Washington Post (2015). Espionage in cyberspace is often a precursor to more aggressive actions, providing adversaries with critical knowledge to plan attacks or influence decision-making processes.
Another significant goal is the disruption of critical infrastructure. Cyber operations targeting power grids, water supplies, or transportation systems can create widespread chaos and compromise public safety. The 2015 cyberattack on Ukraine’s power grid, which left hundreds of thousands without electricity Cybersecurity and Infrastructure Security Agency (2016), illustrates how adversaries can leverage cyber capabilities to undermine a nation’s operational stability. Such attacks often aim to weaken a target’s resilience, demonstrating vulnerability and eroding public trust in essential services.
Cyber warfare is also used to destabilize the economy. Adversaries can disrupt financial institutions, manipulate stock markets, or launch ransomware attacks to cripple economic activities. The WannaCry ransomware attack in 2017, which affected hundreds of organizations worldwide, including hospitals, disrupted critical operations and caused financial losses Gregory (2022). Such actions can destabilize national economies, weaken political systems, and provide adversaries with strategic and monetary gains.
@@ 2.3 Domains
While cyberspace serves as the primary domain for cyber warfare, its effects are far-reaching, often intersecting with the physical domains of land, sea, air, and space. Cyber actions can have tangible impacts in these areas, creating real-world consequences that extend beyond digital boundaries. For instance, cyberattacks targeting the electrical grid can disrupt power supply to entire regions, affecting households, businesses, and critical services. Similarly, satellite hacking can compromise GPS systems, affecting aviation, maritime navigation, and military operations. A notable example is the alleged 2014 cyberattack on a German steel mill, where hackers accessed industrial control systems, causing significant physical damage to a blast furnace BBC News (2014). These intersections highlight the potential for cyber operations to inflict harm in the physical world, blurring the line between cyber and traditional warfare.
Critical infrastructure is another high-stakes target in cyber warfare, given its importance for social functioning and security. Power plants, water supply systems, transportation networks, and healthcare facilities are increasingly reliant on interconnected digital systems, making them vulnerable to cyberattacks. For example, the 2021 Colonial Pipeline ransomware attack disrupted fuel supplies across the southeastern United States, causing widespread panic and economic impact Beerman et al. (2023). Similarly, attacks on healthcare networks, such as the ransomware incident affecting Ireland’s health service in 2021, delayed medical treatments and jeopardized patient safety Reuters Staff (2021). These incidents underscore the critical need for robust cybersecurity measures to protect infrastructure that underpins everyday life and national security.
The human domain is also a key intersection, as cyber warfare increasingly exploits social vulnerabilities to achieve its objectives. Psychological operations (psyops), misinformation campaigns, and social engineering tactics are used to manipulate individuals or groups Yang et al. (2025b). For example, misinformation campaigns during elections aim to influence public opinion, undermine trust in democratic institutions, or polarize societies. In the realm of social engineering, phishing attacks trick individuals into divulging sensitive information, enabling further cyber intrusions. The 2016 US presidential election saw allegations of state-sponsored misinformation campaigns on social media platforms, demonstrating how human perceptions and behaviors can be influenced through cyber means Yang et al. (2025d). By targeting the human domain, adversaries amplify the impact of their cyber operations, turning individuals into unwitting participants in their strategies.
@@ 2.4 Resources
AI resources are increasingly integral to modern cyber warfare, enabling automated tasks that enhance operational efficiency and scalability Rjoub et al. (2023); Mohamed (2025). Artificial intelligence can be used for real-time reconnaissance, identifying vulnerabilities in adversary networks more quickly than human operators. It is also pivotal in crafting sophisticated misinformation campaigns, such as generating deepfake videos to undermine public trust. In addition, AI-driven tools can execute cyberattacks autonomously, selecting targets or adapting strategies based on changing conditions. For example, AI-powered bots might be deployed in a DDoS attack, overwhelming targeted servers with traffic while dynamically adjusting patterns to evade countermeasures.
@@ 2.5 Operational Elements
At the highest level, strategy encompasses comprehensive planning that guides the sequencing of operations to accomplish long-term objectives. For example, a nation-state may strategically prioritize disabling an adversary’s critical infrastructure during a conflict to create chaos and weaken its operational capabilities. This type of strategic planning often involves identifying key targets, allocating resources, and coordinating multiple campaigns to maximize impact and efficiency.
Operations are practical implementations of strategies that encompass both defensive and offensive efforts in cyber warfare. Defensive operations focus on protecting systems by mitigating vulnerabilities, detecting threats, and neutralizing attacks. For instance, a government agency might deploy advanced intrusion detection systems to thwart cyber intrusions targeting sensitive data. These measures are essential to maintain the integrity and security of critical infrastructure. In contrast, offensive operations involve proactive measures aimed at disrupting or disabling adversary systems. An example of this would be launching a cyberattack against an enemy’s command-and-control servers to disrupt their military operations. Such actions are designed to impair the adversary’s capabilities and gain a strategic advantage.
To effectively understand and execute these operations, the kill chain concept has been introduced. Similar to traditional military operations, the kill chain in cyberspace involves understanding how adversaries plan and execute cyber attacks. This framework can be used to capture and model offensive operations, allowing defenders to anticipate and counteract potential threats. Similarly, a defense chain can be utilized to describe defensive operations. This concept outlines the steps taken to protect systems from cyber threats, emphasizing the importance of a structured approach to defense in the ever-evolving landscape of cyber warfare. Together, these frameworks provide valuable insight into offensive and defensive strategies in cyberspace.
@@
Within operational frameworks, information operations play a critical role. These involve tactics such as spreading misinformation, employing deception, and conducting psychological operations (psyops) to manipulate perceptions and gain tactical advantages Roberts and Venables (2024). For example, during a geopolitical crisis, a state might disseminate false information to create confusion and delay the adversary’s response, setting the stage for a decisive offensive cyber operation.
Attacks represent the specific engagements within cyber warfare, akin to battles in traditional conflict. These can take various forms. For example, Advanced Persistent Threats (APTs) are a class of sophisticated, stealthy intrusions that persist over extended periods, such as an APT group infiltrating a financial system to siphon funds or gather intelligence Alshamrani et al. (2019). DDoS Attacks can overwhelm a network with traffic to render it unusable, often employed to disrupt critical services like online banking or government portals. In addition, Jamming Communications blocks or interferes with communication signals, potentially crippling military coordination or emergency response systems.
Finally, tactics involve the specific methods and techniques used to execute these attacks. These are often guided by established frameworks like MITRE ATT&CK Strom et al. (2018), which provides detailed documentation of adversarial Techniques, Tactics, and Procedures (TTPs). For example, a cyber-criminal group might use a phishing campaign to deliver malware (technique), followed by privilege escalation (tactic), to gain control over critical systems. By studying and leveraging frameworks such as MITRE ATT&CK, defenders can better anticipate and counter these tactics, enhancing their resilience against evolving threats.
@@ 2.6 External Factors
Cyber operations are significantly shaped by external considerations, which establish the boundaries and constraints within which cyber actors operate. Among these factors, compliance with legal and regulatory frameworks is paramount. Nations, organizations, and alliances must adhere to international treaties, national laws, and organizational policies that govern the conduct of cyber operations. For example, the Tallinn Manual provides guidelines on the applicability of international law to cyber warfare, defining acceptable conduct during conflicts Schmitt (2017). Similarly, organizations must comply with standards such as the General Data Protection Regulation (GDPR) in the European Union, which governs the handling of personal data and restricts certain cyber activities. Non-compliance can result in legal penalties, diplomatic fallout, or reputational damage, emphasizing the need for operations that align with established norms.
Constraints also play a critical role in shaping cyber operations, encompassing technical, geopolitical, and resource limitations. Technological constraints, such as the availability of advanced tools or expertise, may restrict the scope and sophistication of an operation. For instance, a state with limited cybersecurity infrastructure might struggle to defend against or execute APTs. Geopolitical factors, such as alliances or rivalries, influence the planning and execution of cyber campaigns. A nation might hesitate to launch a cyberattack if it risks escalating tensions with powerful adversaries or alienating allies. Resource limitations, including financial constraints or human resource shortages, can further hinder operations. An example is a small nation relying on third-party contractors for cybersecurity defense, which may leave it vulnerable to more sophisticated state-sponsored attackers.
Finally, ethics introduces an important dimension to cyber operations, dictating what actions are considered acceptable based on societal norms or organizational principles. Ethical considerations often address questions such as whether deception, misinformation, or harm to civilian systems is permissible. For instance, targeting a hospital’s network during a conflict would be deemed unethical and potentially a violation of international humanitarian law, even if the attack disrupts an adversary’s operations. On the other hand, cyber deception Zhu et al. (2025), such as deploying honeypots to mislead attackers, is often considered an ethical and effective defensive measure. However, ethical norms can vary widely across cultures and organizations, leading to differing interpretations of acceptable conduct. For example, some nations may see cyber espionage as a legitimate extension of traditional intelligence-gathering, while others view it as a violation of sovereignty.
@@ 2.7 Outcomes
The outcomes of cyber warfare are multifaceted and can be evaluated at various levels, depending on the objectives and the broader context of the conflict. At the most basic level, operations are assessed as wins, losses, or neutrals, reflecting their immediate effectiveness. For example, a successful DDoS attack that temporarily disrupts an adversary’s critical infrastructure might be considered a “win” in the context of the operation, achieving its intended goal of creating disruption. Conversely, if a cyber defense system thwarts such an attack without significant impact, it would count as a “loss” for the attacker. Neutral outcomes occur when the operation fails to produce meaningful effects or when the benefits and costs balance out, leaving neither party significantly advantaged.
However, the outcomes of cyber warfare extend beyond isolated engagements, encompassing multilevel outcomes that reflect the interplay between individual operations and overarching strategic goals. In this context, success or failure in a specific operation, a “battle”, does not necessarily determine the success or failure of the larger campaign, the “war.” For instance, an APT operation might fail to infiltrate a high-value target due to robust defenses. Although this may seem like a loss, the reconnaissance conducted during the attempt could yield valuable intelligence that enables future operations to succeed, contributing to a longer-term strategic objective.
@@
It is important to recognize that outcomes in cyber warfare are not always zero-sum. Both parties can suffer losses, and the consequences can extend beyond the primary actors, resulting in significant collateral damage. Cyber warfare can be characterized using a framework defined by key elements. Each element corresponds to a game-theoretic counterpart. Clearly defining these elements facilitates the formulation of a game and provides valuable insights for its analysis, which will be explored in the next section.
@@ 3.1 Multi-Echelon Warfare
At the policy level, cyber warfare is linked to grand strategy, also known as high strategy, which extends beyond the immediate concerns of military and cyber operations. Grand strategy integrates economic relations, diplomatic behavior, and international collaboration to achieve long-term objectives. It involves leveraging all available instruments of national power, diplomatic, informational, military, and economic (the DIME framework) to ensure the security and prosperity of a nation or coalition. This level addresses the nation-state scale, where policies shape the operational and tactical choices that follow.
@@
At the strategy level, the focus is on high-level planning to align operations with the overarching objectives of warfare. This involves developing a sequence of defensive and offensive operations, allocating resources to maximize effectiveness, and ensuring coordination between interdependent operations. For example, if the objective is to disrupt an adversary’s critical infrastructure, the strategy could involve a combination of information operations to create misinformation and psychological pressure, followed by a cyber offensive targeting the operating technology systems of a power grid. In this context, information operations, such as the spread of false information about the grid’s stability, could create confusion and panic, thereby softening defenses and aiding the subsequent offensive operation. Resource allocation plays a pivotal role at this level, ensuring that skilled personnel, zero-day vulnerabilities, and AI-powered tools are optimally deployed in planned operations.
@@
At the operational level, the focus narrows to planning specific operations that align with strategic objectives. This involves designing a sequence of attacks that, when executed, achieve the intended outcome of an operation. For instance, in an offensive operation aimed at disabling a transportation system, the plan might involve multiple steps: first, a phishing campaign to gain initial access to the transportation agency’s network; next, a ransomware attack to encrypt critical operational data and disable system functionality; and finally, a DoS attack to prevent recovery efforts. These attacks are not isolated; they are composed so that each amplifies the impact of the others, ensuring the operation achieves its intended goal. The success of an operation depends on the ability to identify vulnerabilities, sequence attacks effectively, and adapt to the adversary’s countermeasures.
@@
The tactic level focuses on the detailed execution of individual attacks (defenses), determining the specific techniques, tactics, and procedures (TTPs) required to carry them out successfully. For example, in a phishing attack, tactics might include crafting highly personalized emails that mimic trusted sources and leveraging psychological manipulation to increase the likelihood that the target will click malicious links. In a DoS attack, tactics might involve deploying botnets to overwhelm servers with traffic while simultaneously masking the source of the attack to delay mitigation. The MITRE ATT&CK framework is often used at this level to plan and refine tactics, ensuring that each attack is optimized for maximum impact.
@@
The three levels of cyber warfare are highly interconnected. Strategic planning provides the overarching framework that dictates the goals and resource allocation for operations. The operational level translates these strategic goals into actionable plans composed of multiple coordinated attacks (defenses). The tactical level ensures that each individual attack (defense) within an operation is meticulously planned and executed to effectively contribute to the overall success. For example, a campaign to disrupt an adversary’s communication infrastructure might strategically plan to exploit interdependencies between networks and users, operationally design a series of cascading attacks on routers and satellites, and tactically deploy precision malware to exploit specific vulnerabilities in routers.